A company has collected data for both sales and commissions over a twelve month period. The average number of sales over three month period per branch is 87 sales per month with a standard deviation of 5 sales per month. The average commission was $5225 per sale with a standard deviation of $773 per month.

a.) Calculate the Coefficient of Variation for both sales and commissions.

b.) Based on the results of part a above which of sales or commissions appears to have had more inherent variability, that is, is inherently more unstable, and why?
Answer #1

a) Coefficient of Variation for sales =(std deviation/mean)=5/87=0.0575 ~ 5.75%

and Coefficient of Variation for commissions =(std deviation/mean)=773/5225=0.1479 ~ 14.79%

b)

as Coefficient of Variation is higher for commission ; therefore it is more unstable or have more inherent variability